what is the limit of cosx-2sinx over cosx+2sinx as x approaches to π/2.
1
Expert's answer
2020-06-08T20:14:09-0400

"\lim\limits_{x\to\frac{\pi}{2}}\frac{\cos x-2\sin x}{\cos x+2\sin x}=[\frac{0-2}{0+2}]=-1"